Problem #1 7r/12=-14
Problem #2 1/3 s + 7/9 = 4/3 s
Use the appropriate property os eqaulity to solve each problem.

Answer by stanbon(75887)

Problem #1
7r/12=-14
Divide both sides by 7 to get:
r/12 = -2
Multiply both sides by 12 to get:
r = -24
==================
Problem #2
1/3 s + 7/9 = 4/3 s
Multiply thru by 9 to get:
3s + 7 = 12s
Subtract 3s from both sides to get:
9x = 7
Divide both sides by 9 to get:
x = 7/9
================
Cheers,
Stan H.